• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name gilman

    Log entry time 16:07:35 on July18,2004

    Entry number 128872

    keyword=day shift summary


    Run 2939 running as we come in - note typo in owl shift summary.
    Plan for day: production data. Problem for now is that the
    Hall C charge asymmetry is large, ~100 ppm, rather than < 10 ppm
    as it should be.

    08:14 run 2940 started
    09:11 run 2941 started
    10:09 run 2942 started
    The Compton DAQ has crashed again. I restart it. The auto-
    analysis of the last run appeared to fail. In the new run,
    it appears to only rarely lock, so I guess we are mostly
    measuring background.
    11:10: 0.52 % bleedthrough measurment, half wave plate inserted
    11:10 run 2943 started
    12:06 run 2944 started
    cryotarget ioc rebooted at ~12:30
    13:03 run 2945 started
    14:00 run 2946 started - flipper inadvertantly run just before run started
    ***septum lost - right septum tripped***
    Trip occurred about 1:00, no alarm noticed. No big jumps in
    beam current. Lumi problem Astrid noticed may have been partly
    due to this.

    while recovering septum, rfzone problem and no bea, for rest of shift.
    kent ans lisa take over to fix up charge asymmetry and recover septum
    during this time.